OpenStudy (ttop0816):

help??? root1-sin^2x/tan^2x+1?? is it cos^2x??

Parth (parthkohli):

\[\dfrac{\sqrt{1 - \sin^2 x}}{\tan^2 x + 1} = \dfrac{\sqrt{\cos^2x}}{\sec^2 x} = \dfrac{\cos x}{\sec^2 x}\]

Parth (parthkohli):

Well, actually, \(\sqrt{\cos^2 x} = |\cos x|\)... just assumed cos(x) for simplicity.

Parth (parthkohli):

Now, can you complete it?

OpenStudy (anonymous):

that right but u can divide an get tan(x/2)=1

Parth (parthkohli):

Hmm?

OpenStudy (ttop0816):

then it would be 1?... these are the answer choices! 0 1 cos^2θ cot^2θ

OpenStudy (anonymous):

yea 1

Parth (parthkohli):

Wait, is that \(\sqrt{1 - \sin^2x}\) or \(1 - \sin^2x\)?

OpenStudy (ttop0816):

its squared both numerator & denominator!

Parth (parthkohli):

Is this the question?\[\dfrac{1 - \sin^2 x}{\tan^2x + 1}\]

OpenStudy (ttop0816):

with the squareroot covering the whole thing at once!

Parth (parthkohli):

Ah! Sorry, lol.

Parth (parthkohli):

\[\dfrac{\cos^2 x}{\sec^2 x } = \cos^4 x\]The root of that is?

OpenStudy (ttop0816):

cos^2x?

Parth (parthkohli):

That's it!

OpenStudy (ttop0816):

yay! then i was correct afterall! (: